OpenStudy (psymon):

Find an equation of a plane that contains all the points that are equidistant from the given points: (-5,1.-3), (2,-1,6) Not even sure how to go about it. Given solution is: 7x-2y+9z = 3

OpenStudy (anonymous):

midpoint and normal

OpenStudy (psymon):

@pgpilot326 That doesn't seem to get me anywhere near the given solution O.o

OpenStudy (anonymous):

there's a line going through the 2 points, right? hlaf way between is the midpoint. a plane that is normal to the line will be equidistant from both given points

OpenStudy (psymon):

Well, the answer I got doing that gave me something kinda, sorta close, just I guess there must've been a specific method that resulted in the solution given. The solution I'm given is x-z = 0. Using the midpoint and the normal vector, I come up with x+ 8y + z = 0

OpenStudy (psymon):

Hmm....hang on, lemme recheck that, I think I made a mistake.

OpenStudy (psymon):

Hmm...nvm, I'm just too unsure xD I simply found the normal vector by taking the cross-product of the two points.

OpenStudy (psymon):

Grr, damn it, the solution I posted was wrong, that was the solution for a different question @_@

OpenStudy (psymon):

Brain cramping today -_-

OpenStudy (anonymous):

midpoint = <(-5+2)/2, (-1+1)/2 , (-3+6)/2> = (-3/2, 0 , 3/2)

OpenStudy (psymon):

Yeah, got that.

OpenStudy (psymon):

And used the cross-product of the two points to get <3, 24, 3>

OpenStudy (psymon):

Im just trying to see if I can find the correct given solution.

OpenStudy (anonymous):

why cross at this point?

OpenStudy (psymon):

I thought you said I need the normal?

OpenStudy (anonymous):

yeah, but what are you crossing?

OpenStudy (psymon):

The inital points given.

OpenStudy (anonymous):

how will that work?

OpenStudy (psymon):

Dunno. I wasn't sure what else to cross to find the normal vector.

OpenStudy (anonymous):